OpenStudy (anonymous):

if AC = 2, CD = 6, and CE = 3, what is BC? .. 12, 11 , 9 or 4

OpenStudy (mathmath333):

is their any diagram given

OpenStudy (anonymous):

Yes

OpenStudy (mathmath333):

so attach the diagram

OpenStudy (anonymous):

/Users/eileenquinones/Desktop/Screen Shot 2014-12-10 at 5.12.48 PM.png

OpenStudy (mathmath333):



OpenStudy (mathmath333):

in \(\triangle ACB \) and \(\triangle ECD\) 1.) \(\angle ACB\cong \angle ECD\)-----(vertical angle) as lines AB and DE are parallel hence 2.) \(\angle ABC\cong \angle EDC\)------------(alternate interior angle) 3.) \(\angle BAC\cong \angle CED\)------------(alternate interior angle) so \(\triangle ACB \) \(\approx\) \(\triangle ECD\) by \(AAA \) similarity theorm hence \(\dfrac{AC}{BC}=\dfrac{CE}{CD}\) \(\dfrac{2}{BC}=\dfrac{3}{6}\) \(BC=\dfrac{2\times 6}{3}\\ ~\\~\\~\\~ \\~\\ \huge BC=4\)

OpenStudy (anonymous):

thank you!